OpenStudy (anonymous):

pleas help me! How does 2√27 = 6√6 ?

OpenStudy (phi):

It doesn't

OpenStudy (raden):

LHS not equal by RHS

OpenStudy (anonymous):

really? its in my math book .. ? O.o

OpenStudy (phi):

you are the victim of a typo. They meant \[ 2 \sqrt{27}= 2 \sqrt{9\cdot 3} =2 \cdot 3 \sqrt{3} =6 \sqrt{3} \]

OpenStudy (anonymous):

yes good job phi

OpenStudy (anonymous):

oh, thank you! that makes so much more sense!